# PENEX: AdaBoost-Inspired Neural Network Regularization

### v2

@@ Abstract
AdaBoost sequentially fits so-called weak learners to minimize an exponential loss, which penalizes mislabeled data points more severely than other loss functions like cross-entropy. Paradoxically, AdaBoost generalizes well in practice as the number of weak learners grows. In the present work, we introduce Penalized Exponential Loss (PENEX), a new formulation of the multi-class exponential loss that is theoretically grounded and, in contrast to the existing formulation, amenable to optimization via first-order methods. We demonstrate both empirically and theoretically that PENEX implicitly maximizes margins of data points. Also, we show that gradient increments on PENEX implicitly parameterize weak learners in the boosting framework. Across computer vision and language tasks, we show that PENEX exhibits a regularizing effect often better than established methods with similar computational cost. Our results highlight PENEX’s potential as an AdaBoost-inspired alternative for effective training and fine-tuning of deep neural networks.
@@ 1 Introduction
A seemingly unrelated branch of literature to regularization is boosting (Schapire, 1990): The method of generating a strong learner by sequentially fitting weak learners. A weak learner can be thought of as an inaccurate “rule-of-thumb” that barely does better than random guessing. A prime example of a weak learner is a single-split decision tree (“decision stump”), which classifies based on a single attribute split (e.g., Hastie et al. (2017, p. 362); left part of Fig. 1). The goal of boosting is to train weak learners in order to gradually refine the predictions arising from a linear combination over all weak learners that have been fit so far (resulting in a single, strong learner). One of the best-known boosting algorithms for classification is AdaBoost (Freund & Schapire, 1995), which has been called the “best off-the-shelf classifier in the world” by Leo Breiman (Friedman et al., 2000). Although originally derived from the probably approximately correct learning framework (Valiant, 1984), Breiman (1999) famously gave AdaBoost an interpretation in the language of empirical risk minimization (Vapnik, 2000) by showing that AdaBoost minimizes an exponential loss function. This loss has been criticized for its lack of robustness with respect to outliers and misclassified data points (Bishop, 2006, p. 662). Counterintuitively, empirical studies have demonstrated that AdaBoost resists “overfitting” in practice as the number of weak learners increases (Drucker & Cortes, 1995; Breiman, 1996a; Friedman et al., 2000), thereby seemingly defying the Occam’s razor principle by which simpler models should be favored over complex ones (e.g., MacKay (2003, p. 343)). To resolve the paradox, various authors have developed theoretical arguments, most of them based on margin maximization (Schapire et al., 1998; Rätsch et al., 2001; Rosset et al., 2003; Grove & Schuurmans, 1998; Onoda et al., 1998; Rosset et al., 2004): [^1] Though alternative theories based on bias-variance theory (Breiman, 1996a, b; Bauer & Kohavi, 1999) and self-averaging properties (Wyner et al., 2017) exist. The margin is the minimum distance of a data point to any decision boundary. Due to the large penalty on miss-classifications, this margin is maximized and thus, the model concentrates more on hard-to-learn examples, which play a similar role to support vectors (Cortes & Vapnik, 1995). This principle is demonstrated in Fig. 2. Maximizing the margin for many training data points, in turn, guarantees good generalization (e.g., Schapire et al. (1998)).
@@
Motivated by the works above, we take the view that AdaBoost does not generalize well in spite of the exponential loss, but because of it, in the sense that it acts as a regularizer. The main contribution of the present work is to effectively translate this regularizing quality to training deep neural network classifiers. Specifically, we introduce the Penalized Exponential Loss (PENEX), which reformulates the objective of the multi-class exponential loss (Zhu et al., 2009) to a neural network-friendly variant by replacing a hard sum constraint by a SumExp penalty. We show that PENEX is Fisher consistent (Fisher, 1922), i.e., it is a proper loss for estimating the Bayes-optimal classifier. We also prove formally that PENEX effectively maximizes margins with high probability, thereby providing a theoretical foundation for PENEX’s effectivity. We propose to minimize PENEX via gradient descent, which, at first sight, appears to contrast the idea of sequentially fitting weak learners in the boosting framework. However, we show that small gradient-based parameter increments implicitly parameterize a form of weak learner, as visualized in Fig. 1. This result draws an analogy in that training on PENEX keeps improving on generalization as the model is trained for many epochs, just as how AdaBoost keeps improving on generalization as many decision stumps are added.
@@
Empirically, we observe that PENEX often improves upon the performance of popular regularizers such as label smoothing (Szegedy et al., 2016) and confidence penalty (Pereyra et al., 2017) across computer vision and language modeling tasks.
@@
Figure 2: Comparison of Margins. Neural networks trained with PENEX (center) tend to implicitly maximize margins (here, geometric margins, indicated for an example point in green), in a similar way to support vector machines (right), here trained with a RBF kernel (Schölkopf & Smola, 2002, p. 46). Training neural networks via cross-entropy loss (left), in contrast, typically leads to smaller margins.
@@
  • We introduce PENEX (Sec. 2), an unconstrained alternative to the optimization problem proposed by Zhu et al. (2009) as stated in Sec. 2 below. We demonstrate favorable theoretical properties of PENEX, specifically Fisher consistency (Prop. 2.1) and its margin maximization quality (Thm. 2.1).
  • We demonstrate the relationship between performing gradient descent on PENEX and AdaBoost. Specifically, we can interpret the former as an implicit form of AdaBoost, where small parameter increments correspond to parameters of weak learners (Fig. 1; Sec. 2.3).
  • We empirically compare the regularizing effect of PENEX with ablations and other regularizers, on diverse tasks ranging from computer vision to fine-tuning of a language model. Our results show that PENEX typically outperforms existing regularization techniques with similar computational cost, in low data regimes (Sec. 4).
@@ 2 Penalized Exponential Loss (PENEX)
For $f:\mathcal{X}\mapsto\mathbb{R}^{K}$ mapping from input space $\mathcal{X}$ to $K$ classes, the exponential loss (EX) is given as
@@
The intuition behind the constraint in Sec. 2 is that it acts as a barrier preventing the $f^{(j)}$ from diverging without bounds. In the present work, we propose an alternative to CONEX, which we refer to as penalized exponential loss (PENEX). It penalizes EX 1 by a SumExp over $f^{(j)}$, specifically
@@
is the SumExp and $\rho>0$ is a penalty parameter. The penalty in 3 fulfills a similar purpose as the zero-sum constraint in Sec. 2, but acts differently: PENEX penalizes individual logits for being large rather than requiring logits to “cancel each other out”. The main benefit of PENEX 3 over CONEX Sec. 2 is that PENEX avoids constraints, which would require more elaborate algorithms like the augmented Lagrangian method (Hestenes, 1969). Instead, 3 poses just an objective function, which can be optimized directly using gradient-based algorithms like Adam (Kingma & Ba, 2015).
@@ 2.1 Theoretical Properties
Proposition 2.1 (Fisher Consistency).
For any $\rho>0$ and $\alpha>0$, the penalized exponential loss 3 is Fisher consistent, meaning that the minimizer of the population equivalent [^2] This means that we exchange $\hat{\mathbb{E}}$ by $\mathbb{E}$ in 3., called $f_{*}$, recovers the Bayes-optimal classifier in the sense that
$P(y\;|\;\mathbf{x})\;\propto\;\exp\left\{\left(1+\alpha\right)f_{*}^{(y)}(\mathbf{x})\right\},$
i.e., the $f_{*}^{(y)}(\mathbf{x})$ recover the true logits (up to constant shift) at temperature $(1+\alpha)^{-1}$. [^3] Based on this result, in practice, we always rescale logits during inference (see Sec. C.1).
A proof for Prop. 2.1 is supplied in Sec. A.1. While other loss functions such as cross-entropy $\mathcal{L}_{\mathrm{CE}}(f)\coloneqq-\hat{\mathbb{E}}\big[\log\big(\hat{P}(y\,|\,\mathbf{x})\big)\big]$ with $\hat{P}(y\,|\,\mathbf{x})\propto\exp\{f^{(y)}(\mathbf{x})\}$ are Fisher consistent as well, the solutions on finite datasets typically differ, as can be seen in Fig. 3: When the “correct” logit grows negative ($f^{(1)}(\mathbf{x})\rightarrow$-$\infty$), the gradient of PENEX keeps decreasing. The cross‐entropy gradient, however, levels out at 1-1.
Popular regularization methods for deep learning are $\ell_{2}$ regularization, confidence penalty and label smoothing. These methods can be interpreted as adding a regularization term to cross-entropy (Meister et al., 2020) and, in contrast to PENEX, fail Fisher consistency:
Observation 2.1.
Consider the family of loss functions obtained by adding a regularization term $\Omega(f)$ to the cross-entropy loss:
$\mathcal{L}_{\mathrm{reg}}(f;\,\lambda)\;=\;\mathcal{L}_{\mathrm{CE}}(f)\;+\;\lambda\,\Omega(f).$ (4)
In general, $\mathcal{L}_{\mathrm{reg}}(f;\,\lambda)$ is not Fisher consistent for $\hat{P}(y\,|\,\mathbf{x})\propto\exp\{f^{(y)}(\mathbf{x})\}$.
We demonstrate a proof of Obs. 2.1 in Sec. A.2. A key aspect in the generalization ability of PENEX lies in an implicit margin-maximization property. In the present work, we define the margin $m(\mathbf{x},y)$ of an example $(\mathbf{x},y)$ as the minimum difference of the true logit to any “false” logit, that is
$m_{f}(\mathbf{x},y)\;\coloneqq\;f^{(y)}(\mathbf{x})\,-\,\max_{j\neq y}f^{(j)}(\mathbf{x}).$ (5)
The next theorem shows that PENEX effectively aims at maximizing the margin for many data points:
Theorem 2.1 (Margin Maximization).
@@
$\mathbb{P}(m_{f}(\mathbf{x},y)\leq\gamma)\leq e^{\frac{\gamma\alpha}{\alpha+1}}\rho^{-\frac{\alpha}{\alpha+1}}\mathbb{E}[\mathcal{L}_{\mathrm{\scriptscriptstyle PENEX}}(f;\,\alpha,\rho)].$ (6)
A proof is provided in Sec. A.4. Intuitively, Thm. 2.1 says that the probability of a small margin for a data point is low, if PENEX has a small value for a model $f$, in expectation. Not only does Thm. 2.1 explain the margin-maximization property of PENEX (Fig. 2), but also it gives us a hint on how to reasonably choose $\rho$ as a function of $\alpha$:
@@
The parameter $\rho$ that minimizes the upper bound in 6 is
$\rho_{*}(\alpha,f)\;=\;\alpha\frac{\mathbb{E}[\mathcal{L}_{\mathrm{\scriptscriptstyle EX}}(f;\alpha)]}{\mathbb{E}[\mathrm{SE}(f;\mathbf{x})]}.$ (7)
This result follows from realizing that the upper bound in 6 is convex and differentiable in $\rho$, thus we can solve for $\rho_{*}$ by setting its derivative to zero.
2.2 Practical Implementation
In practice, the optimal penalty parameter in 7 cannot be computed directly, so we rely on batch statistics for the estimation. To further reduce variance, we use exponential moving averaging (EMA) and clipping, as shown in Alg. 1. The weights for EMA and clipping are fixed for all experiments, i.e., we do not treat them as tuning parameters. We note that Alg. 1 can be implemented in a stateful loss function object and does not require a custom training loop. We refer to App. C for more details regarding the implementation.

### v3

@@ Abstract
AdaBoost sequentially fits so-called weak learners to minimize an exponential loss, which penalizes misclassified data points more severely than other loss functions like cross-entropy. Paradoxically, AdaBoost generalizes well in practice as the number of weak learners grows. In the present work, we introduce Penalized Exponential Loss (PENEX), a new formulation of the multi-class exponential loss that is theoretically grounded and, in contrast to the existing formulation, amenable to optimization via first-order methods, making it a practical objective for training neural networks. We demonstrate that PENEX effectively increases margins of data points, which can be translated into a generalization bound. Empirically, across computer vision and language tasks, PENEX improves neural network generalization in low-data regimes, often matching or outperforming established regularizers at comparable computational cost. Our results highlight the potential of the exponential loss beyond its application in AdaBoost.
Exponential Loss
@@ 1 Introduction
A seemingly unrelated branch of literature to regularization is boosting (Schapire, 1990): The method of generating a strong learner by sequentially fitting weak learners. A weak learner can be thought of as an inaccurate “rule-of-thumb” that barely does better than random guessing. A prime example of a weak learner is a single-split decision tree (“decision stump”), which classifies based on a single attribute split (e.g., Hastie et al. (2017, p. 362); left part of Fig. 1). The goal of boosting is to train weak learners in order to gradually refine the predictions arising from a linear combination over all weak learners that have been fit so far (resulting in a single, strong learner). One of the best-known boosting algorithms for classification is AdaBoost (Freund & Schapire, 1995), which has been called the “best off-the-shelf classifier in the world” by Leo Breiman (Friedman et al., 2000). Although originally derived from the probably approximately correct learning framework (Valiant, 1984), Breiman (1999) famously gave AdaBoost an interpretation in the language of empirical risk minimization (Vapnik, 2000) by showing that AdaBoost minimizes an exponential loss function. This loss has been criticized for its lack of robustness with respect to outliers and misclassified data points (Bishop, 2006, p. 662). Counterintuitively, empirical studies have demonstrated that AdaBoost resists “overfitting” in practice as the number of weak learners increases (Drucker & Cortes, 1995; Breiman, 1996a; Friedman et al., 2000), thereby seemingly defying the Occam’s razor principle by which simpler models should be favored over complex ones (e.g., MacKay (2003, p. 343)). To resolve the paradox, various authors have developed theoretical arguments, most of them based on margin properties (Schapire et al., 1998; Rätsch et al., 2001; Rosset et al., 2003; Grove & Schuurmans, 1998; Onoda et al., 1998; Rosset et al., 2004): [^1] Though alternative theories based on bias-variance theory (Breiman, 1996a, b; Bauer & Kohavi, 1999) and self-averaging properties (Wyner et al., 2017) exist. The margin is the minimum distance of a data point to any decision boundary. Due to the large penalty on miss-classifications, the model concentrates more on hard-to-learn examples, which play a similar role to support vectors (Cortes & Vapnik (1995), Fig. 2). Thus, the tail of the margin distribution is minimized. A light tail of the margin distribution, in turn, guarantees good generalization (Mohri et al., 2018, Thm. 9.2).
@@
Motivated by the works above, we take the view that AdaBoost does not generalize well in spite of the exponential loss, but because of it, in the sense that it acts as a regularizer. The main contribution of the present work is to effectively translate this regularizing quality to training deep neural network classifiers. Specifically, we introduce the Penalized Exponential Loss (PENEX), which reformulates the objective of the multi-class exponential loss (Zhu et al., 2009) to a neural network-friendly variant by replacing a hard sum constraint by a SumExp penalty. We show that PENEX is self-calibrating, meaning that at the population level its optimal logits can be rescaled in closed form to recover the true class-probability distribution, and that it is Fisher consistent (Fisher, 1922), i.e., its optimal logits induce the Bayes-optimal classifier (Lin, 2004; Zhu et al., 2009). We also prove that PENEX bounds the tail of the margin distribution, thereby providing a theoretical foundation for PENEX’s generalization ability. [^2] We include a broader discussion about how this result differs from the known result by Schapire et al. (1998) in App. B. We propose to minimize PENEX via gradient descent, which, at first sight, appears to contrast the idea of sequentially fitting weak learners in the boosting framework. However, we show that small gradient-based parameter increments implicitly parameterize a form of weak learner, as visualized in Fig. 1. This result draws an analogy in that training on PENEX keeps improving on generalization as the model is trained for many epochs, just as how AdaBoost keeps improving on generalization as many decision stumps are added.
@@
Empirically, in low data regimes, we observe that PENEX often improves upon the performance of popular regularizers such as label smoothing (Szegedy et al., 2016) and confidence penalty (Pereyra et al., 2017) across computer vision and language modeling tasks.
@@
Figure 2: Comparison of Margins. Neural networks trained with PENEX (center) tend to implicitly increase the smallest margins (here, geometric margins, indicated for an example point in green), in a similar way to support vector machines (right), here trained with a RBF kernel (Schölkopf & Smola, 2002, p. 46). Training via cross-entropy loss (left), in contrast, typically leads to smaller margins.
@@
  • We introduce PENEX (Sec. 2), an unconstrained alternative to the optimization problem proposed by Zhu et al. (2009) as stated in Sec. 2 below. We demonstrate favorable theoretical properties of PENEX, specifically self-calibration (Prop. 2.1), Fisher consistency (Cor. 2.1) and its margin-promoting quality (Thm. 2.1). The latter can be combined with an existing result by Mohri et al. (2018) to provide a generalization bound (Cor. 2.2).
  • We demonstrate the relationship between performing gradient descent on PENEX and AdaBoost (Prop. 2.3). Specifically, we can interpret the former as an implicit form of AdaBoost, where small parameter increments correspond to parameters of weak learners (Fig. 1; Sec. 2.3).
  • We empirically compare the regularizing effect of PENEX with ablations and other regularizers, on diverse tasks ranging from computer vision to fine-tuning of a language model. Our results show that PENEX often improves upon existing regularization techniques with similar computational cost, in low data regimes (Sec. 4).
@@ 2 Penalized Exponential Loss (PENEX)
For class logits $f:\mathcal{X}\mapsto\mathbb{R}^{K}$ mapping from input space $\mathcal{X}$ to $K$ classes, the exponential loss (EX) is given as
@@
The intuition behind the constraint in Sec. 2 is that it acts as a barrier preventing the $f^{(j)}$ from diverging without bounds: without the constraint, the exponential loss 1 can be decreased indefinitely by sending all logits to infinity. In the present work, we propose an alternative to CONEX, which we refer to as penalized exponential loss (PENEX). It penalizes EX 1 by a SumExp over $f^{(j)}$, specifically
@@
is the SumExp and $\rho>0$ is a penalty parameter. The penalty in 3 fulfills a similar purpose as the zero-sum constraint in Sec. 2, but acts differently: PENEX penalizes individual logits for being large rather than requiring logits to “cancel each other out”. The penalty is an alternative means of making the exponential loss 1 well-posed. The main benefit of PENEX 3 over CONEX Sec. 2 is that PENEX avoids constraints, which would require more elaborate algorithms like the augmented Lagrangian method (Hestenes, 1969). Instead, 3 poses just an objective function, which can be optimized directly using gradient-based algorithms like Adam (Kingma & Ba, 2015).
@@ 2.1 Theoretical Properties
Proposition 2.1 (Self-Calibration).
For any $\rho>0$ and $\alpha>0$, and for any $\mathbf{x}$ in the support of the marginal distribution of $\mathbf{x}$, the minimizer of the population equivalent of 3, called $f_{*}$:
$f_{*}\coloneqq\arg\min_{f}\,\mathbb{E}[\mathcal{L}_{\mathrm{\scriptscriptstyle PENEX}}(f;\,\alpha,\rho)],$ (4)
is unique and recovers the true class probabilities $P(y\mid\mathbf{x})$ via
$P(y\mid\mathbf{x})\;\propto\;\exp\!\left\{(1+\alpha)\,f_{*}^{(y)}(\mathbf{x})\right\},$
i.e., the $f_{*}^{(y)}(\mathbf{x})$ recover the log-posteriors at temperature $(1+\alpha)^{-1}$, up to a constant shift.
A proof for Prop. 2.1 is supplied in Sec. A.1. Prop. 2.1 shows that PENEX provides a closed-form temperature re-scaling to recover the true class probabilities, on the population level. In addition, Prop. 2.1 lays the ground for showing that PENEX is Fisher consistent:
Corollary 2.1 (Fisher Consistency).
Under the assumptions of Prop. 2.1, PENEX 3 is Fisher consistent according to the definition of Zhu et al. (2009), i.e., $f_{*}$ 4 induces the Bayes-optimal classifier via
$\arg\max_{i}P(i\mid\mathbf{x})\;=\;\arg\max_{i}f_{*}^{(i)}(\mathbf{x}).$
Cor. 2.1 follows directly from Prop. 2.1 since $(1+\alpha)>0$ and so
$\arg\max_{i}f_{*}^{(i)}(\mathbf{x})$ $=\arg\max_{i}\exp\bigl\{(1+\alpha)\,f_{*}^{(i)}(\mathbf{x})\bigr\}$
$=\arg\max_{i}P(i\mid\mathbf{x})\,.$
While other loss functions such as cross-entropy $\mathcal{L}_{\mathrm{CE}}(f)\coloneqq-\hat{\mathbb{E}}\big[\log\big(\hat{P}(y\,|\,\mathbf{x})\big)\big]$ with $\hat{P}(y\,|\,\mathbf{x})\propto\exp\{f^{(y)}(\mathbf{x})\}$ are Fisher consistent as well, the solutions on finite datasets usually differ, as can be inferred from Fig. 3: When the “correct” logit grows negative ($f^{(1)}(\mathbf{x})\rightarrow$-$\infty$), the gradient of PENEX keeps decreasing. The cross‐entropy gradient, however, levels out at 1-1.
A key aspect in the generalization ability of PENEX lies in a margin-emphasizing quality. In the present work, we define the margin $m(\mathbf{x},y)$ of an example $(\mathbf{x},y)$ as the minimum difference of the true logit to any “false” logit,
$m_{f}(\mathbf{x},y)\;\coloneqq\;f^{(y)}(\mathbf{x})\,-\,\max_{j\neq y}f^{(j)}(\mathbf{x}).$
The next theorem shows that PENEX effectively aims at avoiding small margins:
Theorem 2.1 (Margin Tail Bound).
@@
$\hat{\mathbb{P}}(m_{f}(\mathbf{x},y)\leq\gamma)\leq e^{\gamma\frac{\alpha}{\alpha+1}}\rho^{-\frac{\alpha}{\alpha+1}}\mathcal{L}_{\mathrm{\scriptscriptstyle PENEX}}(f;\,\alpha,\rho),$ (5)
where $\hat{\mathbb{P}}$ denotes empirical probability.
A proof is provided in Sec. A.3. Intuitively, Thm. 2.1 says that the fraction of data points with a small margin is low, if PENEX has a small value for a model $f$. Defining the expected miss-classification error $\mathcal{R}$ as
$\mathcal{R}(f)\;\coloneqq\;\mathbb{E}[\mathbbm{1}\{\mathrm{arg}\,\mathrm{max}_{i}\,f^{(i)}(\mathbf{x})\,\neq\,y\}],$ (6)
we can combine Thm. 2.1 with a known result from statistical learning (Mohri et al., 2018, Thm. 9.2) to obtain the following generalization bound:
Corollary 2.2 (Generalization Bound).
Assume $f\in\mathcal{F}$ and denote Rademacher complexity of $\mathcal{F}$ by $\mathfrak{R}$. Then, for any $\delta\in(0;1),\rho>0,\gamma>0$, it holds with probability at least $1-\delta$ that
$\mathcal{R}(f)\;\leq\;e^{\gamma\frac{\alpha}{\alpha+1}}\rho^{-\frac{\alpha}{\alpha+1}}\mathcal{L}_{\mathrm{\scriptscriptstyle PENEX}}(f;\,\alpha,\rho)\,+$
$\frac{4K}{\gamma}\mathfrak{R}(\mathcal{F})\,+\,\sqrt{\frac{\log(1/\delta)}{2n}}.$
We interpret Cor. 2.2 by minimizing the right-hand side over $\gamma>0$. The choice of $\gamma$ mediates a trade-off: the capacity term $\frac{4K}{\gamma}\mathfrak{R}(\mathcal{F})$ decreases as $1/\gamma$, whereas the leading term increases as $\exp(\gamma\tfrac{\alpha}{\alpha+1})$. Consequently, the bound suggests an intermediate margin threshold $\gamma_{\ast}$. Notably, achieving a smaller PENEX value decreases the prefactor of the exponential term, which shifts $\gamma_{\ast}$ toward larger values and thereby reduces the capacity term $\frac{4K}{\gamma_{\ast}}\mathfrak{R}(\mathcal{F})$.
In addition to providing a generalization bound, Thm. 2.1 suggests a principled choice of $\rho$:
@@
The parameter $\rho$ that minimizes the upper bound in 5 is
$\rho_{*}(\alpha,f)\;=\;\alpha\frac{\hat{\mathbb{E}}[\mathcal{L}_{\mathrm{\scriptscriptstyle EX}}(f;\alpha)]}{\hat{\mathbb{E}}[\mathrm{SE}(f;\mathbf{x})]}.$ (7)
This result follows from realizing that the upper bound in 5 is convex and differentiable in $\rho$, thus we can solve for $\rho_{*}$ by setting its derivative to zero.
2.2 Notes on the Practical Implementation
In practice, the optimal penalty parameter in 7 is expensive to compute directly, because $f$ changes during training. Thus, we rely on online batch statistics for estimation. To reduce variance, we use exponential moving averaging (EMA) and clipping, as shown in Alg. 1. The weights for EMA and clipping are fixed for all experiments, i.e., we do not treat them as tuning parameters. We note that Alg. 1 can be implemented in a stateful loss object and does not require a custom training loop. We refer to App. D for more details regarding the implementation.
Also, the practical implementation implements the fixed temperature-rescaling suggested by Prop. 2.1. That is, during inference, estimates of class logits are always multiplied by $1+\alpha$ before feeding them into the softmax function (more details are provided in Sec. D.1).